Operation Summary for January 3, 2011 - January 5, 2011 |
|
Monday Jan. 3 The day shift began with Operations monitoring the accelerator complex while most of the complex was in access during the continuing Tevatron (TEV) sector B1 vacuum work. During the day shift, Operations reported accesses into the TEV A-E enclosures and the Transfer Hall, F-Sector, Main Injector (MI), and into the Antiproton Source (Pbar): Rings, Transport, Pre-Target, and Pre-Vault areas. Later in the day, operators prepared MI-65, MiniBooNE, and Linac for accesses. Safety system tests occurred throughout the day shift. Linac personnel began working on a RF station (LRF3) at 2 PM, replacing its 7651 tube. During the evening shift, the MI work ended at the beginning of shift. Linac personnel completed the repairs on LRF3. Operations established beam to MiniBooNE at 5:55 PM. Operations restored beam to NuMI at 6:25 PM. Tuesday Jan. 4 The midnight shift began with the TEV sector B1 vacuum leak-checking finishing up, and with NuMI and MiniBooNE taking beam. During the midnight shift, Mechanical Support personnel finished their vacuum work at B1. Cryo personnel began preparations for starting the B1 cryo system. During the day shift, Safety System personnel conducted interlock testing, from 8 to 10:30 AM, of F-Sector and the Transfer Hall. Safety personnel re-cored the Pbar service building AP-30 to a different key set due to reverse proton Mu2E beam studies. Operators searched and secured the D0 collision hall. Linac personnel switched the Preaccelerator from using the H- Source to the I- Source. During the evening shift, operators searched and secured the TEV A-E enclosures. Cryo system personnel began the TEV sector B1 cooldown at 7:20 PM. Operations reported that the Pbar extraction Lambertson (D:ELAM) failed during Mu2E studies. Wednesday Jan. 5 The midnight shift began with Operations monitoring the accelerator complex, with MI and Recycler operating normally or conducting studies, and with NuMI and MiniBooNE taking beam. During the midnight shift, operators began preparing TeV A-E power supplies for testing. Operations reported that Safety System personnel began conducting TeV A-E interlock testing at the end of shift. Operations turned off the Pbar Rings for Mu2E reconfiguration of beam detectors and for D:ELAM repairs. The Plans for Wednesday The plans for today are to continue with safety system tests, to monitor the B1 cryo system cooldown, to conduct a TeV hipot, to allow Pbar personnel to conduct studies, and to get back to stashing and stacking by this evening. |
|
More Information Comments and Suggestions |
| last modified 01/05/2011 email Fermilab |
| Security, Privacy, Legal | |